# Accuracy-Robustness Trade Off via Spiking Neural Network Gradient Sparsity Trail

### v1

@@ 4.2 Natural adversarial robustness of SNN
4.3 Theoretical explanation and gradient analysis
@@ 5 How does shared sparsity trail in gradient computation affect performance?
Previous studies have demonstrated a strong relationship between the magnitude of weight gradients and model performance Luu et al. (2024b); Fang et al. (2021a); Shalev-Shwartz et al. (2017). Building on this foundation, we aim to investigate how sparsity in the input gradients may suppress weight gradients during optimization, thereby influencing overall performance. Consider a multilayer SNN of the form:
where is the input signal, are the model parameters at layer and is the hard-reset IF neurons with membrane potential , defined as:
@@
As observed from the derivation results, input gradient shared a gradient chain with the Jacobian of during backward computation and a shorter chain of with respect to , leading to an intuitive indication that:
@@
Parameters that are placed closer to the initial input signal will shared a longer gradient chain, making sparsity in input gradient have a stronger correlation with parameter gradient (a more illustrative demonstration can be found at Figure 4).
So in this scenario, some observations can be obtained as:
Input gradient sparsity positively correlated with null entries in weights and : .
Input gradient sparsity positively correlated with weight gradient sparsity of over null entries of :

### v2

@@ 4.2 Natural adversarial robustness of SNN
4.3 Explanation and analysis
@@
As observed from the theorem, adversarial vulnerability scale positively with non-zero entries in input gradient, we can form a corollary as:
Corollary 1.
The adversarial vulnerability scale positively with non-zero entries of input gradient .
@@ 5 How does shared sparsity trail in gradient computation affect performance?
Previous studies have demonstrated a strong relationship between the magnitude of weight gradients and model performance (Luu et al., 2024b; Fang et al., 2021a; Shalev-Shwartz et al., 2017). Building on this foundation, we aim to investigate how sparsity in the input gradients may suppress weight gradients during optimization, thereby influencing overall performance. To prove this statement, we consider the following theorem:
Theorem 2.
Given a multilayer SNN of the form:
where is the input signal, are the model parameter matrices at layer and is the hard-reset IF neurons with membrane potential , defined as:
Then as a result, we would have the following:
1.
2.
Parameters that are placed closer to the initial input signal will shared a longer gradient chain, making sparsity in input gradient have a stronger correlation with parameter gradient.
Proof.
@@
As obtained from the derivation results, input gradient shared a gradient chain with the Jacobian of during backward computation and a shorter chain of with respect to , leading us to several observations:
@@
Parameters that are placed closer to the initial input signal will shared a longer gradient chain, making sparsity in input gradient have a stronger correlation with parameter gradient111A more illustrative demonstration can be found at Figure 4
We are confident that any sparsity in the shared gradient path propagates backward to both the weight and input gradients. Still, opposite of sparsity, how can we formally quantify the relationship between the density of the weight gradient (specifically of , as it is being the closest parameter to input and is highly affected by the shared gradient trail) with respect to input gradient, particularly in the context of Theorem 2? To address this, we establish the following theorem, which provides a precise characterization of their relationship:
Theorem 3.
Given the same condition as Theorem 2, the parameter matrix at layer is said to have a row/column density degree of / (or /-dense row/column for short) if each of its row/column vector (where , ) has at most non-zero entries , defined as:
and for a vector as:
If the norm of input gradient and weight gradient can be defined as:
then norm of weight gradient of can bounded as:
Proof.
Each entry in the input gradient is defined as:
and the probability of an element in the sum chain being dense is upper bounded as:
As long as there is an element in the sum chain is non-zero, the entry of gradient is non-zero. So, from here we can quantify the bound of the probability of input gradient entry being dense as:
and the expectation of the input gradient norm across matrix is:
Similar derivation process can be applied to the weight gradient and obtain its expectation. We start with an element of the sum chain as:
Divide obtained inequalities from Equation 16 and 17:
As observed from derivation results, weight gradient density is bounded by input gradient density by a factor of . ∎